# Reservoir Predictive Path Integral Control for Unknown Nonlinear Dynamics

### v1

@@ V Numerical Experiments
    QPMPC [23]: the input is obtained by minimizing the cost in 4. Because the ESN is nonlinear, the model is linearized via a first-order Taylor expansion at every control cycle to yield a QP formulation.
@@
    MPPI: the input is computed by Algorithm 3 with ${\tilde{K}}=1$ and $\tilde{\Sigma}=0$.
@@
    UMPPI: the input is computed by Algorithm 3 with ${\tilde{K}}\geq 1$ and $\tilde{\Sigma}>0$.
The hyperparameters are determined through preliminary experiments prior to the execution of Algorithm 1. To verify the concept of simultaneous online identification and control, different plant parameters are used during hyperparameter search and the main experiment. The ESN hyperparameters, namely the sparseness and spectral radius of ${\mathbf{W}}^{\text{res}}$, the leak rate $\alpha$, and the range of the input weight matrix ${\mathbf{W}}^{\text{in}}$, are determined through random search to minimize the cost function 8 when random input signals are applied. Furthermore, the MPPI hyperparameter, the inverse temperature $\lambda$, is determined through random search to minimize the cost function (4) when Algorithm 1 is executed using the ESN model with the optimal hyperparameters.
All numerical computations in this study were performed on a host machine running Windows 11 (CPU: Intel Core i9-13900@2.00 GHz, 24 cores; memory: 64 GB) under WSL2 (Ubuntu 22.04). The numerical computation code was implemented in Julia 1.10. In QPMPC, QP problems were formulated using the JuMP package [36] and solved with the OSQP package [37]. In RPPI and URPPI, 16-thread parallelization on the CPU was employed to perform parallel sampling of the input sequences. All source code and data needed to reproduce the results are publicly available at https://github.com/ToyotaCRDL/rppi.
@@ V-A Duffing Oscillator
The experimental parameters for the Duffing oscillator system are specified below. During hyperparameter tuning, the Duffing oscillator was configured with linear stiffness $k=1.0$, nonlinear stiffness $k_{\text{nl}}=1.0$, mass $m=1.0$, and damping $c=1.0$. For the main experiments these values were changed to $k=1.1$, $k_{\text{nl}}=0.9$, $m=1.1$, and $c=0.9$. The control period was fixed at $\Delta s=0.1$, and 38 was integrated with an explicit Euler scheme using this time step. Cost function weighting employs $Q=100$ for output tracking and $R=1$ for input regulation. The reference trajectory alternates between positions $\boldsymbol{y}^{\text{ref}}=1.0$ and $\boldsymbol{y}^{\text{ref}}=-1.0$ every $20$ seconds. Control inputs are constrained to the admissible set ${\mathcal{U}}=[-10,10]^{1\times H}$. The ESN parameters are configured as follows. The state matrix uses a dimension of $\hat{N}=400$ with sparsity factor $\rho=0.48$ (ratio of non-zero connections) and spectral radius of $0.89$. Input weights ${\mathbf{W}}^{\text{in}}$ are set to uniform random values in the range [2.0,2.0][-2.0,2.0]. The network employs a leak rate of $\alpha=0.59$ and the hyperbolic tangent activation function $f(x)=\tanh(x)$. For RLS implementation, output weights ${\mathbf{W}}_{-1}^{\mathrm{out}}$ are initialized as uniform random values in [0.1,0.1][-0.1,0.1], the precision matrix is set to $^{\mathbf{P}}_{-1}=I_{\hat{N}}$, and the discount factor is $\gamma=1.0$. The control parameters are specified as follows. MPPI operates with the input sample size $K=1000$, prediction horizon $H=10$, inverse temperature $\lambda=1.9$, and noise covariance $\Sigma=1.9$. UMPPI extends these parameters with the output sample size ${\tilde{K}}=10$ and output weight covariance $\tilde{\Sigma}=1.0$.
@@
Figure 2 presents the time evolution of system variables under each control method. Figure 2a displays the output response ${\mathbf{y}}_{t}$, while Fig. 2b shows the corresponding control input ${\mathbf{u}}_{t}$. Given the stochastic nature of ESN initialization, MPPI sampling, and UMPPI perturbations, experimental results exhibit inherent variability across executions. Consequently, experiments were conducted across 20 random seeds; for each seed, we independently resampled (i) the ESN state matrix and input-weight matrix, (ii) the initial output-weight matrix, (iii) the MPPI input-perturbation sequences, and (iv) the UMPPI output-weight perturbation sequences. Ensemble means are shown as solid lines, and standard deviations are depicted as shaded regions. The results demonstrate that while QPMPC fails to achieve asymptotic convergence to the reference trajectory, both MPPI and UMPPI successfully attain asymptotic tracking performance. Moreover, QPMPC exhibits substantially higher variability in both control signals and system outputs compared to the reduced variability observed with MPPI and UMPPI.
@@
where $T$ is the terminal time step, here $T=400$. Both MPPI and UMPPI achieve substantial reductions in control cost compared to QPMPC, with UMPPI achieving the lowest average cost among the three. Specifically, UMPPI reduces control cost by approximately 63% compared to QPMPC and by 7% compared to MPPI. In terms of computation time, MPPI requires 0.034 seconds on average to compute control inputs, while UMPPI requires 0.046 seconds. Considering both control cost and computation time, UMPPI provides improved control performance over MPPI at the expense of only a modest increase in computation time.
Figure 4: Averaged continuous-wavelet scalogram of the control signals for the Duffing oscillator. Each signal is analyzed with a complex Morlet wavelet (frequency-domain variance $2\pi$). The horizontal axis shows time, the vertical axis shows frequency, and color indicates the coefficient magnitude. From top to bottom: results for UMPPI, MPPI, and QPMPC.
Figure 4 shows the continuous wavelet scalograms of the control inputs generated by each method. The analysis is confined to the first $6$ seconds of each input signal, with the frequency content at each step visualized. Both MPPI and UMPPI exhibit frequency components distributed over a considerably wider range than QPMPC, which is advantageous for capturing the dynamics of the system being identified. This occurs because MPPI, as a random-sampling-based control, inherently produces input signals containing a broad spectrum of frequencies. Moreover, around the area of $t<3~$\mathrm{s}$\$, the UMPPI input signal displays an even wider spread of frequency components compared to MPPI. UMPPI achieves this by injecting perturbations into the control input based on the identification accuracy of its output weight matrix. This behavior suggests an intrinsic mechanism within UMPPI that automatically balances exploration and exploitation.
@@ V-B Four-Tank System
are shown in Fig. 6. Here, the terminal time step $T$ is set to $T=1000$. Both MPPI and UMPPI achieve substantial reductions in control cost compared to QPMPC, with UMPPI achieving the best overall performance. Specifically, UMPPI reduces control cost by 33.1% compared to QPMPC, and by 3.9% relative to MPPI. In terms of computation time, MPPI requires 0.037 seconds on average to compute control inputs, while UMPPI requires 0.049 seconds. Overall, UMPPI provides improved control performance over MPPI with only a small additional computational cost.
Figure 7: Averaged continuous-wavelet scalogram of the control signals for the four-tank system. Each signal is analyzed with a complex Morlet wavelet (frequency-domain variance $2\pi$). The horizontal axis shows time, the vertical axis shows frequency, and color indicates the coefficient magnitude. From top to bottom: results for UMPPI, MPPI, and QPMPC.
Figure 7 presents the averaged continuous wavelet scalograms of control inputs for UMPPI, MPPI, and QPMPC. We focus on the first $60$ seconds of each trial, with the magnitude of the coefficients visualized at every step. Both MPPI and UMPPI exhibit broader frequency distributions than QPMPC, and, around $t=40~$\mathrm{s}$\$, UMPPI displays the widest spread. These results confirm that MPPI’s random-sampling approach introduces diverse spectral content, and UMPPI’s perturbations further enhance automatic exploration-exploitation balancing.
@@ Acknowledgments
    where ${\mathbf{B}}_{t}\coloneqq\sum_{\tau=0}^{t}\gamma^{t-\tau}{\mathbf{x}}_{\tau}{\mathbf{y}}_{\tau}^{\top}$. From 11 and $4$,
    $\begin{aligned} {\mathbf{W}}^{\mathrm{out}*}_{t}-\bar{\mathbf{W}}^{\mathrm{out}}&={\mathbf{P}}_{t}\left({\mathbf{B}}_{t}-{\mathbf{A}}_{t}\bar{\mathbf{W}}^{\mathrm{out}}\right)\\ &={\mathbf{P}}_{t}\sum_{\tau=0}^{t}\gamma^{t-\tau}{\mathbf{x}}_{\tau}\mbox{\boldmath$\epsilon$}_{\tau}^{\top}.\end{aligned}$ (45)
@@
    where $\epsilon_{\tau,\ell}$ is the $\ell$-th component of $\mbox{\boldmath$\epsilon$}_{\tau}$. Using the fact that
@@
    $\begin{aligned} &{\mathbb{E}}[\mathrm{vec}({\mathbf{W}}^{\mathrm{out}*}_{t}-\bar{\mathbf{W}}^{\mathrm{out}})\mathrm{vec}({\mathbf{W}}^{\mathrm{out}*}_{t}-\bar{\mathbf{W}}^{\mathrm{out}})^{\top}]\\ &=\begin{bmatrix}\sigma_{1}^{2}{\mathbf{P}}_{t}&0&\cdots&0\\ 0&\sigma_{2}^{2}{\mathbf{P}}_{t}&&\vdots\\ \vdots&&\ddots&0\\ 0&\cdots&0&\sigma^{2}_{L}{\mathbf{P}}_{t}\end{bmatrix}\\ &=\tilde{\Sigma}\otimes{\mathbf{P}}_{t},\end{aligned}$ (48)
@@
    For the second term, by taking the reference probability density function $q$ as 25 and 26, it follows
@@
    Substituting LABEL:eq:KL into 49 completes the proof. ∎
  • Proof of Proposition IV.5
    In 49, taking the probability density function of $V_{t}$ and $W_{t}$ as 28, we obtain
@@
    In the third line, $p(V_{t}|U_{t},\Sigma)$ defined in 18 was substituted. When ${\mathcal{U}}={\mathbb{R}}^{M\times H}$, the solution to this optimization problem satisfies ${\mathbf{u}}_{\tau}^{*}={\mathbb{E}}_{p^{*}}[{\mathbf{v}}_{\tau}],\ (\tau=t,\ldots,t+H-1)$.
    Next, by introducing importance sampling using $\hat{U}_{t}$, we show 32. The above optimal input is expressed as
@@
    Using the reference density function $q(V_{t})$ in 25, we have
@@
    Substituting 59 and 61 into 55 completes the proof. ∎

### v2

@@ V Numerical Experiments
    QPMPC [23] (baseline): the input is obtained by minimizing the cost in 3. Because the ESN is nonlinear, the model is linearized via a first-order Taylor expansion at every control cycle to yield a QP formulation.
@@
    MPPI (proposed RPPI): the input is computed by Algorithm 3 with ${\tilde{K}}=1$ and $\tilde{\Sigma}=0$.
@@
    UMPPI (proposed UMPPI): the input is computed by Algorithm 3 with ${\tilde{K}}\geq 1$ and $\tilde{\Sigma}>0$.
The hyperparameters are determined through preliminary experiments prior to the execution of Algorithm 1. To verify the concept of simultaneous online identification and control, different plant parameters are used during hyperparameter search and the main experiment. The ESN hyperparameters, namely the sparsity and spectral radius of ${\mathbf{W}}^{\text{res}}$, the leak rate $\alpha$, and the range of the input weight matrix ${\mathbf{W}}^{\text{in}}$, are determined through random search to minimize the cost function 10 when random input signals are applied. Furthermore, the MPPI hyperparameter, the inverse temperature $\lambda$, is determined through random search to minimize the cost function (3) when Algorithm 1 is executed using the ESN model with the optimal hyperparameters.
All numerical simulations were performed on a workstation equipped with an Intel Core i7-12700K processor (12 cores, 20 threads) and 24 GB RAM, running Ubuntu 24.04.1 LTS on Windows Subsystem for Linux 2 (WSL2). The numerical computation code was implemented in Julia 1.10. In QPMPC, QP problems were solved with the OSQP package [42]. At each time step, the ESN model is linearized around the current estimated state $\hat{\mathbf{x}}_{t}$ to formulate the QP problem. In RPPI and URPPI, 16-thread parallelization on the CPU was employed to perform parallel sampling of the input sequences. All source code and data needed to reproduce the results are publicly available at https://github.com/ToyotaCRDL/rppi.
@@ V-A Duffing Oscillator
The experimental parameters for the Duffing oscillator system are specified below. During hyperparameter tuning, the Duffing oscillator was configured with linear stiffness $k=1.0$, nonlinear stiffness $k_{\text{nl}}=1.0$, mass $m=1.0$, and damping $c=1.0$. For the main experiments these values were changed to $k=1.1$, $k_{\text{nl}}=0.9$, $m=1.1$, and $c=0.9$. The control period was fixed at $\Delta s=0.1$, and 39 was integrated with an explicit Euler scheme using this time step. Cost function weighting employs $Q=100$ for output tracking and $R=1$ for input regulation. The reference trajectory alternates between positions $\boldsymbol{y}^{\text{ref}}=1.0$ and $\boldsymbol{y}^{\text{ref}}=-1.0$ every $20$ seconds. Control inputs are constrained to the admissible set ${\mathcal{U}}=[-10,10]^{1\times H}$. The ESN parameters are configured as follows. The state matrix uses a dimension of $\hat{N}=400$ with sparsity factor $\rho=0.48$ (ratio of non-zero connections) and spectral radius of $0.89$. Input weights ${\mathbf{W}}^{\text{in}}$ are set to uniform random values in the range [2.0,2.0][-2.0,2.0]. The network employs a leak rate of $\alpha=0.59$ and the hyperbolic tangent activation function $f(x)=\tanh(x)$. For RLS implementation, output weights ${\mathbf{W}}_{-1}^{\text{out}}$ are initialized as uniform random values in [0.1,0.1][-0.1,0.1], the precision matrix is set to $^{\mathbf{P}}_{-1}=I_{\hat{N}}$, and the discount factor is $\gamma=1.0$. The control parameters are specified as follows. MPPI operates with the input sample size $K=1000$, prediction horizon $H=10$, inverse temperature $\lambda=1.9$, and noise covariance $\Sigma=1.9$. UMPPI extends these parameters with the output sample size ${\tilde{K}}=10$ and output weight covariance $\tilde{\Sigma}=1.0$.
@@
Figure 2 presents the time evolution of system variables under each control method. Figure 2a displays the output response ${\mathbf{y}}_{t}$, while Fig. 2b shows the corresponding control input ${\mathbf{u}}_{t}$. Given the stochastic nature of ESN initialization, MPPI sampling, and UMPPI perturbations, experimental results exhibit inherent variability across executions. Consequently, experiments were conducted across 20 random seeds; for each seed, we independently resampled (i) the ESN state matrix and input-weight matrix, (ii) the initial output-weight matrix, (iii) the MPPI input-perturbation sequences, and (iv) the UMPPI output-weight perturbation sequences. Ensemble means are shown as solid lines, and standard deviations are depicted as shaded regions. The results demonstrate that while QPMPC fails to achieve asymptotic convergence to the reference trajectory, both MPPI and UMPPI successfully attain asymptotic tracking performance. Moreover, QPMPC exhibits higher variability in both control signals and system outputs compared to the reduced variability observed with MPPI and UMPPI.
@@
TABLE I: Comparison of average control cost and computation time among control methods on the Duffing oscillator.
Method Control Cost Computation Time $\mathrm{s}\text{/}$]
QPMPC 3151 0.008
MPPI (Proposed) 1247 0.030
UMPPI (Proposed) 1155 0.044
@@
where $T$ is the terminal time step, here $T=400$. Both MPPI and UMPPI achieve substantial reductions in control cost compared to QPMPC, with UMPPI achieving the lowest average cost among the three. Specifically, UMPPI reduces control cost by approximately 63% compared to QPMPC and by 7% compared to MPPI. This performance improvement can be attributed to UMPPI’s ability to account for model uncertainty. By minimizing the expected cost over the distribution of ESN parameters, UMPPI generates control inputs that are robust to identification errors, avoiding aggressive actions that might be optimal for a nominal model but risky under uncertainty. As shown in Table I, MPPI requires 0.030 seconds on average to compute control inputs, while UMPPI requires 0.044 seconds. Although UMPPI incurs a slight computational overhead due to the additional sampling of output weights, it remains sufficiently fast for online control with a control period of 0.1 seconds. Considering both control cost and computation time, UMPPI provides improved control performance over MPPI at the expense of only a modest increase in computation time.
@@ V-B Four-Tank System
TABLE II: Comparison of average control cost and computation time among control methods on the four-tank system.
Method Control Cost Computation Time $\mathrm{s}\text{/}$]
QPMPC 2088 0.008
MPPI (Proposed) 1451 0.033
UMPPI (Proposed) 1395 0.053
@@
are shown in Fig. 5. Here, the terminal time step $T$ is set to $T=1000$. Both MPPI and UMPPI achieve substantial reductions in control cost compared to QPMPC, with UMPPI achieving the best overall performance. Specifically, UMPPI reduces control cost by 33.1% compared to QPMPC, and by 3.9% relative to MPPI. Similar to the Duffing oscillator case, the stochastic formulation of UMPPI allows it to cope with the uncertainty in the learned four-tank dynamics, leading to more robust control performance. As shown in Table II, MPPI requires 0.033 seconds on average to compute control inputs, while UMPPI requires 0.053 seconds. Although UMPPI incurs a slight computational overhead due to the additional sampling of output weights, it remains sufficiently fast for online control with a control period of 1.0 seconds. Overall, UMPPI provides improved control performance over MPPI with only a small additional computational cost.
@@ Acknowledgments
where ${\mathbf{B}}_{t}\coloneqq\sum_{\tau=0}^{t}\gamma^{t-\tau}{\mathbf{x}}_{\tau}{\mathbf{y}}_{\tau}^{\top}$. From 13 and 45,
$\begin{aligned} {\mathbf{W}}^{\text{out}*}_{t}-\bar{\mathbf{W}}^{\text{out}}&={\mathbf{P}}_{t}\left({\mathbf{B}}_{t}-{\mathbf{A}}_{t}\bar{\mathbf{W}}^{\text{out}}\right)\\ &={\mathbf{P}}_{t}\sum_{\tau=0}^{t}\gamma^{t-\tau}{\mathbf{x}}_{\tau}\mbox{$\epsilon$}_{\tau}^{\top}.\end{aligned}$ (46)
@@
where $\epsilon_{\tau,\ell}$ is the $\ell$-th component of $\mbox{$\epsilon$}_{\tau}$. Using the fact that
@@
$\begin{aligned} &{\mathbb{E}}[\mathrm{vec}({\mathbf{W}}^{\text{out}*}_{t}-\bar{\mathbf{W}}^{\text{out}})\mathrm{vec}({\mathbf{W}}^{\text{out}*}_{t}-\bar{\mathbf{W}}^{\text{out}})^{\top}]\\ &=\begin{bmatrix}\sigma_{1}^{2}{\mathbf{P}}_{t}&0&\cdots&0\\ 0&\sigma_{2}^{2}{\mathbf{P}}_{t}&&\vdots\\ \vdots&&\ddots&0\\ 0&\cdots&0&\sigma^{2}_{L}{\mathbf{P}}_{t}\end{bmatrix}\\ &=\tilde{\Sigma}\otimes{\mathbf{P}}_{t},\end{aligned}$ (49)
@@
For the second term, by taking the reference probability density function $q$ as 26 and 27, it follows
@@
Substituting LABEL:eq:KL into 50 completes the proof. ∎
In 50, taking the probability density function of $V_{t}$ and $W_{t}$ as 29, we obtain
@@
In the third line, $p(V_{t}|U_{t},\Sigma)$ defined in 19 was substituted. When ${\mathcal{U}}={\mathbb{R}}^{M\times H}$, the solution to this optimization problem satisfies ${\mathbf{u}}_{\tau}^{*}={\mathbb{E}}_{p^{*}}[{\mathbf{v}}_{\tau}],\ (\tau=t,\ldots,t+H-1)$.
Next, by introducing importance sampling using $\hat{U}_{t}$, we show 33. The above optimal input is expressed as
@@
Using the reference density function $q(V_{t})$ in 26, we have
@@
Substituting 60 and 62 into 56 completes the proof. ∎